I am asked to simply this problem:
6^3a+2 multiplied by 6^2-2a over 6^a+5
(6 to the 3a+2 power multiplied by 6 to the 2nd power minus 2a power over 6 to the A power plus 5th power
6^3a+2 multiplied by 6^2-2a
---------------------------
6^a+5
Thank you

Answer by sdmmadam@yahoo.com(530)

6^3a+2 multiplied by 6^2-2a
---------------------------
6^a+5
=[(6)^(3a+2)]X[(6)^(2-2a)]X[1/(6)^(a+5)] (using (ab)/c = (ab)X(1/c) )
={(6)^[(3a+2)+(2-2a)]}X[1/(6)^(a+5)]
(when quantities of the same base are multiplied
write the base and add the powers)
=(6)^[3a-2a+2+2]X(6)^[-(a+5)] ( Using the rule (1/a^m) = (a)^(-m))
=(6)^(a+4) X (6)[-(a+5)]
=(6)^[(a+4)-(a+5)]
(when quantities of the same base are multiplied
write the base and add the powers)
=(6)^(a+4-a-5)
=(6)^(4-5)
=(6)^(-1)
=1/(6)^1
=1/6